OpenStudy (daisy.xoxo619):

G: A PARALLEL B <1=X+3Y <2=2X+30 <3=5Y+20 FIND M<1 WHAT I KNOW : 2X+30+5Y+20=180

OpenStudy (daisy.xoxo619):



OpenStudy (mathstudent55):

Where is angle 1?

OpenStudy (daisy.xoxo619):



OpenStudy (mathstudent55):

Ok. What you wrote above is correct. Angles 2 and 3 are same side interior angles, so they are supplementary. That means the sum of their measures is 180. That gives you one equation. What are angles 1 and 3 called?

OpenStudy (daisy.xoxo619):

CORRESPONDING THERE CONG

OpenStudy (mathstudent55):

Can you please use small case letters?

OpenStudy (mathstudent55):

Yes, that's correct. What do you know about corresponding angles of parallel lines?

OpenStudy (mathstudent55):

Oh, I see, you already wrote cong. Ok, angles 1 and 3 are congruent. What equation do you get from that?

OpenStudy (daisy.xoxo619):

x+3y=5y+20

OpenStudy (mathstudent55):

Great. Now you have two equations in two unknowns. You can solve the system of equations for x and y. Here is the system of equations: \(2x+30+5y+20=180\) \(x+3y=5y+20\) Do you know how to solve a system of equations?

OpenStudy (daisy.xoxo619):

does x=5y-17

OpenStudy (mathstudent55):

No. You need to subtract 3y from both sides in the second equation, not 3.

OpenStudy (mathstudent55):

2x + 30 + 5y + 20 = 180 x + 3y = 5y + 20 Let's combine like terms in the first equation: 2x + 5y + 50 = 180 Now let's subtract 50 from both sides: 2x + 5y = 130 Now, let's subtract 3y to the right in the second equation: x = 2y + 20

OpenStudy (mathstudent55):

The two equations in our system now are: 2x + 5y = 130 x = 2y + 20 Since the second equation is solved for x, use the substitution method and substitute what x is equal to in the first equation. Then solve for x.

OpenStudy (daisy.xoxo619):

does m<1=70

OpenStudy (mathstudent55):

Correct.

OpenStudy (daisy.xoxo619):

thanks

OpenStudy (mathstudent55):

You're welcome.